rearranging equations
make x the subjects of:
x-y=w
x+3y=5v
v2+x=w
5v+x=w
x2=6y
x2+v=w
x2/7=w-y
x2+6/w=y
x/y+7=v
v/5x=w/y
vx-w=6
5w2+vx=y
a(x+b)=c or x=something/something
6(x+a)=b or x=something/something
a(x-5)=b or x=something/something
c(6+x)=5 or x=something/something
I want both solutions for the last 4 questions.

Answer by dabanfield(803)

1.)a(x+b)=c or x=something/something
2.)6(x+a)=b or x=something/something
3.)a(x-5)=b or x=something/something
4.)c(6+x)=5 or x=something/something
I want both solutions for the last 4 questions.
Using the distributive property:
1.) a(x+b) = c
ax+ab = c
ax = c-ab
x = (c-ab)/a
2.) 6(x-a) = b
6x -6a = b
6x = b+6a
x = (b+6a)/6
The last two are similar.